OpenStudy (anonymous):

I'm unclear on how to solve this problem, can anyone help me please? (: "Given the sequence -3,0,3, which of the following is the equivalent to

OpenStudy (anonymous):

\[\sum_{n=5}^{8} a _{n}\]

OpenStudy (anonymous):

54 102 144 162 These are my choices.

OpenStudy (unklerhaukus):

\[\sum_{n=5}^{8} a _{n}=a_5+a_6+a_7+a_8\]

OpenStudy (unklerhaukus):

what is \(-3,0,3,\) ? are they \(a_1, a_2,a_3\) ?

OpenStudy (anonymous):

Im not really sure, I think you're supposed to derive a "summation" from those. When I did that I got 6+3n but I'm not sure if that's right

OpenStudy (anonymous):

I got that from substituting 9 as the first term and 3 as i in the equation \[a_{n}\] = a1 + (n-1)d

OpenStudy (unklerhaukus):

how do you know 9 is the first term?

OpenStudy (anonymous):

Because n=5 so that means that I start on the first term of the original sequence? so -3,0,3,6,9,12,... so 9 is the 5th term, so I started there. From there I continued 9,12,15,18,21,24,27,30,33... and used 30 as the 8th term. I just sorta guessed on that part...

OpenStudy (unklerhaukus):

oh right, so \[a_5=9,\quad a_6=12, \quad a_7=15, \quad a_8=18\]

OpenStudy (anonymous):

Okay, so if those are the values, then what would you do afterwards?

OpenStudy (unklerhaukus):

just add em

OpenStudy (anonymous):

oh, wow okay. So I guess 9 wasn't the first term haha

OpenStudy (unklerhaukus):

\[\sum_{n=5}^{8} a _{n}=9+12+15+18=\]

OpenStudy (anonymous):

I got 54 correct?

OpenStudy (unklerhaukus):

i think thats right

OpenStudy (anonymous):

Okay, thank you so much! (: